OpenStudy (anonymous):

How do i get started on this problem (5x-2)^1/3=2 help please?

OpenStudy (anonymous):

Alright so your equation is \[(5x-2)^{\frac{ 1 }{ 3 }}=2\] right? This isn't so hard. First of all, let's make things a little more easier by removing the power of the equation. Let's cube each side, like so. \[((5x-2)^{\frac{1}{3}})^{3} = 2^{3}\] That gives us \[(5x-2)^{1}=8\] It's simple from here on. \[5x-2=8\] \[5x=10\] \[x=10/5\] \[x=2\] I hope my answer helped you. -swas

OpenStudy (anonymous):

great, thank you!

OpenStudy (anonymous):

If you liked my answer, please hit the Best Response button and close the question :)